# Towards Implicit Prompt For Text-To-Image Models

### v3

@@ 1 Introduction
Beneath the advancements of these benchmarks, we find there exists a notable weakness: they are restrained on explicit prompts, where target objects are directly clarified in the text. Can T2I models generate images without directly clarifying the target? Will indirect descriptions bring risks to T2I generation? In this paper, we answer these questions with implicit prompts, which do not contain the concrete identity of targets while describing relevant target information. In our pilot experiments, we discover that T2I models indeed possess the capability to generate desired images by implicit prompts. For instance, as shown in Figure 1 (a), an implicit prompt of Eiffel Tower “Located in France, an iconic iron lattice tower, …” can induce DALL-E 3 to generate the correct target as the explicit prompt “Eiffel Tower” achieves. Then, the potential risks brought by implicit prompts arise. For example, in Figure 1, implicit prompts can guide T2I models to break the safety constraint. To our knowledge, existing work leaves T2I models toward implicit prompts unexplored. Therefore, we delve into the general performance and application issues toward implicit prompts.
@@
  • NSFW Issue. We found that some implicit prompts could bypass most prompt safety filters of T2I models for their harmless illusion, but can generate images with NSFW content. From a commercial perspective, Midjourney demonstrated a stronger security coefficient compared to the DALL-E series because it can identify more subtle NSFW implications in implicit prompts and prevent their NSFW generation. What’s more, upon our in-depth analysis of these implicit NSFW prompts, we conclude that compared to normal words, certain stylistic professional terminologies, overly specific descriptions of body parts, and words with ambiguity or multiple meanings pose a greater risk of generating NSFW content.
In this position paper, we advocate for more attention on implicit prompts for the T2I community, especially focusing on our proposed three significant aspects. On one hand, implicit prompts show hopeful prospects and enable more scalable generation. With impressive comprehensive ability, T2I models could generate more vivid images, which benefits human creators in increasing speed and reducing skill requirements. On the other hand, existing policy restrictions may be ineffective in addressing implicit prompts. Thus, the privacy and NSFW content deduced by implicit prompts should raise significant concerns. The solution to mitigate the risk of implicit prompts needs to be exploited. We hope that our research can stimulate research on the capability and impacts of implicit prompts.
@@ 4.3 NSFW Issues
5 Discussion and Conclusion
In this section, we analyze the factors that influence different models’ performances and risks across three aspects.
@@ 5 Discussion and Conclusion
Conclusion. In this position paper, we introduce implicit prompts, a type of prompt that describes a target without directly naming it, to broaden the current scope of the Text-to-Image community. Through a comprehensive benchmark tailored for implicit prompts, our investigation across three critical areas—General Symbols, Celebrity Privacy, and NSFW Issues—evaluates the performance of six contemporary T2I models. Our evaluation not only opens up research avenues to develop models toward implicit prompts but also sheds light on their existing capabilities in accurately interpreting such prompts. Significantly, our findings discover a potential risk in violating celebrity privacy and generating NSFW content via implicit prompts, challenges that current policy constraints do not sufficiently address. This paper calls for a more nuanced development of policy constraints to mitigate these risks, highlighting the need for further research and attention in the T2I community.
@@
Broader Impact. This research illuminates the remarkable capabilities of T2I models to interpret and visualize concepts conveyed through implicit prompts, thereby extending the creative and practical applications of text-to-image generation. This advancement could democratize art creation, enhance educational tools, and foster innovation across fields by enabling a more intuitive interaction with T2I systems, where users can generate detailed visual content without needing explicit or technical descriptions. However, the capabilities of these models also introduce profound ethical considerations. The potential risks for T2I models to inadvertently breach celebrity privacy or generate not-safe-for-work (NSFW) content from seemingly benign prompts should raise concerns. Such instances underscore the imperative for robust ethical frameworks and advanced policy constraints that can comprehend the nuances of implicit prompts. It is crucial that developers and users of T2I technology remain cognizant of these risks, advocating for the responsible use of such powerful tools.
@@ Appendix B Detailed Samples of ImplicitBench
We present Table 5 to show the numerical comparative results of various T2I models in inferring from implicit prompts and successfully generating images related to celebrities across different professions
@@
Models Politician(%) Actor(%) Poet(%) Artist(%) Musician(%) Athlete(%) Singer(%) Writer(%) Writer(%) Scientist(%)

### v4

# Position: Towards Implicit Prompt For Text-To-Image Models

@@ 1 Introduction
Beneath the advancements of these benchmarks, we find there exists a notable weakness: they are restrained on explicit prompts, where target objects are directly clarified in the text. Can T2I models generate images without directly clarifying the target? Will indirect descriptions bring risks to T2I generation? In this paper, we answer these questions with implicit prompts, which do not contain the concrete identity of targets while describing relevant target information.
In our pilot experiments, we discover that T2I models indeed possess the capability to generate desired images by implicit prompts. For instance, as shown in Figure 1 (a), an implicit prompt of Eiffel Tower “Located in France, an iconic iron lattice tower, …” can induce DALL-E 3 to generate the correct target as the explicit prompt “Eiffel Tower” achieves. Then, the potential risks brought by implicit prompts arise. For example, in Figure 1, implicit prompts can guide T2I models to break the safety constraint. To our knowledge, existing work leaves T2I models toward implicit prompts unexplored. Therefore, we delve into the general performance and application issues toward implicit prompts.
@@
  • NSFW Issues. We found that some implicit prompts could bypass most prompt safety filters of T2I models for their harmless illusion, but can generate images with NSFW content. From a commercial perspective, Midjourney demonstrated a stronger security coefficient compared to the DALL-E series because it can identify more subtle NSFW implications in implicit prompts and prevent their NSFW generation. What’s more, upon our in-depth analysis of these implicit NSFW prompts, we conclude that compared to normal words, certain stylistic professional terminologies, overly specific descriptions of body parts, and words with ambiguity or multiple meanings pose a greater risk of generating NSFW content.
This position paper advocates for more attention on implicit prompts for the T2I community, especially focusing on our proposed three significant aspects: General Symbols, Celebrity Privacy and NSFW Issues. On one hand, implicit prompts show hopeful prospects and enable more scalable generation. With impressive comprehensive ability, T2I models could generate more vivid images, which benefits human creators in increasing speed and reducing skill requirements. On the other hand, existing policy restrictions may be ineffective in addressing implicit prompts. Thus, the privacy and NSFW content deduced by implicit prompts should raise significant concerns. The solution to mitigate the risk of implicit prompts needs to be exploited. We hope that our research can stimulate research on the capability and impacts of implicit prompts. Our code and benchmark will be released in website https://github.com/yangyue5114/implicit_prompt.
@@ 4.3 NSFW Issues
5 Result Discussion and Dealing Solution
In this section, we first analyze the factors that influence different models’ performances and risks across three aspects. Then we propose some practical solutions that T2I community can adopt to deal with the non-negligible safety dangers when facing implicit prompts.
5.1 Result Discussion
@@
We also counted the gender distribution of celebrities in our implicit prompt benchmark. Then we calculated the accuracy of our evaluated six T2I models in successfully generating images of male and female celebrities. As Figure 8 shown, the result reveals that these models pretend to expose a slightly high identity of male celebrities compared with females, which is generally aligned well with the discovery in other studies on T2I gender bias (Luccioni et al., 2023; Seshadri et al., 2023). Previous studies speculate that this bias may caused by different quantitative proportions in training data on males and females.
Figure 8: The relationship between the gender of celebrities and accuracy rates toward implicit prompts in terms of celebrity.
@@
5.2 Dealing Solution
After our initial exploration toward the implicit prompt for T2I models, we have realized not only the importance of the implicit prompt but also its potential risks. Therefore, we propose three practical and useful solutions to deal with the safety dangers caused by implicit prompts, to further improve the security coefficient and defense ability of current T2I models and community.
T2I Input Stage: Using implicitBench to train a recaption model for safety. Considering the input stage on T2I input prompts, we can utilize our proposed implicit prompt benchmark to tune a prompt recaption model like (Hao et al., 2024), securing any input prompts in celebrity privacy and NSFW issues.
Figure 10: The framework to train a recaption model for secure user input prompt using our proposed implicitBench.
Since our benchmark includes plenty of pairs of <<< implicit prompt, generated unsafe image >>> over ten thousand, and we have designed a series of evaluation methods to compute a result judging whether images contain unsafe pixels. Thus, as shown in Figure 10, we can take the evaluation results as reward scores to fine-tune a large language model (LLM). After fine-tuning, the trained LLM can adaptively optimize the user’s input prompt, ultimately achieving a high safety score. Through this method, we can transform an unsafe implicit prompt into a harmless version, dealing with its unsafety generation in celebrity and NSFW content.
T2I Output Stage: Enhancing the detection ability of image safety checker by our generated edged-unsafe images. Our experiments demonstrate that existing T2I systems’ safety checkers (Machine Vision & Learning Group LMU, 2022; Qu et al., 2023) can only detect some obviously unsafe images and respond by refusing to display them. But current safety checkers failed to detect images generated by our implicit prompts, because our generated images may be edged-unsafe: contain small areas of gore, slightly disturbing content, or pornography located in corners or the edges of the image, which means the edged-unsafe images generated by our implicit prompts are of tremendous value. Therefore, we can utilize these edged-unsafe images to train more powerful and intelligent image safety checkers, along with current training data. This solution would effectively solve the security vulnerabilities caused by implicit prompts.
Add dynamic prompt safety filters with contextual awareness and a user intent clarification mechanism. Develop dynamic safety filters that adjust their sensitivity based on the context of the prompt. Together with a clarification system that prompts users to provide more context or clarify their intent when an implicit prompt is detected as potentially leading to unsafe content generation. This mechanism benefits in identifying potential NSFW content, privacy violations, or other harmful implications while keeping regular usage of common prompts.
6 Conclusion
@@
In this position paper, we introduce implicit prompts, a type of prompt that describes a target without directly naming it, to broaden the current scope of the Text-to-Image community. Through a comprehensive benchmark tailored for implicit prompts, our investigation across three critical areas—General Symbols, Celebrity Privacy, and NSFW Issues—evaluates the performance of six contemporary T2I models. Our evaluation not only opens up research avenues to develop models toward implicit prompts but also sheds light on their existing capabilities in accurately interpreting such prompts. Significantly, our findings discover a potential risk in violating celebrity privacy and generating NSFW content via implicit prompts, challenges that current policy constraints do not sufficiently address. This paper calls for a more nuanced development of policy constraints to mitigate these risks, highlighting the need for further research and attention in the T2I community.
Impact Statement
@@
This research illuminates the remarkable capabilities of T2I models to interpret and visualize concepts conveyed through implicit prompts, thereby extending the creative and practical applications of text-to-image generation. This advancement could democratize art creation, enhance educational tools, and foster innovation across fields by enabling a more intuitive interaction with T2I systems, where users can generate detailed visual content without needing explicit or technical descriptions. However, the capabilities of these models also introduce profound ethical considerations. The potential risks for T2I models to inadvertently breach celebrity privacy or generate not-safe-for-work (NSFW) content from seemingly benign prompts should raise concerns. Such instances underscore the imperative for robust ethical frameworks and advanced policy constraints that can comprehend the nuances of implicit prompts. It is crucial that developers and users of T2I technology remain cognizant of these risks, advocating for the responsible use of such powerful tools.
Acknowledgements
This work is partially supported by the National Key R&D Program of China (NO.2022ZD0160101, NO.2022ZD0161000), and Science and Technology Commission of Shanghai Municipality (STCSM) (No. 21511101100). This work was done during his internship at Shanghai Artificial Intelligence Laboratory.
@@ Appendix B Detailed Samples of ImplicitBench
We present Table 5 to show the numerical comparative results of various T2I models in inferring from implicit prompts and successfully generating images related to celebrities across different professions.
@@
Models Politician Actor Poet Artist Musician Athlete Singer Writer Writer Scientist